A. Transform the following logarithmic expressions to its exponential form or vice versa. 1. 24 = 16 2. 9² = 81 3. 3-2 1 9 = 4. log39 = 2 5. log164 = ²/ B. Evaluate the following logarithmic expressions. deposition 1. log232 2. log,1 3. log381 4. log16913 1 5. logs /
1
Expert's answer
2021-10-29T00:41:02-0400

A.

1.

"\log_216=4"

2.


"\log_981=2"

3.


"\log_3\dfrac{1}{9}=-2"

4.

"3^2=9"

5.


"16^{1\/2}=4"

B.

1.

"\log_232=5"

2.


"\log1=0"

3.


"\log_381=4"

4.


"\log_{169}13=\dfrac{1}{2}"

5.


"\log_s1=0, s>0, s\not=1"